Solve:
5(x-2)^2 = 3
Thank you for helping me!!!

Answer by stanbon(75887)

Solve:
5(x-2)^2 = 3
5(x^2-4x+4)=3
5x^2-20x+17=0
Use the quadratic formula to get:
x=[20+-sqrt(400-4*5*17)]/10
x=[20+-sqrt60]/10
x=[20+-2sqrt15]/10
x=2+(1/5)sqrt15 or x=2-(1/5)sqrt15
-------------
Cheers,
Stan H.